Circular No. 8001

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

POSSIBLE NOVA IN SAGITTARIUS V. Tabur, Wanniassa, Australian Capital Territory, reports his discov-

ery of an apparent nova (red mag 9.7) on unfiltered CCD images taken with a 100-mm telephoto lens on Oct. 25.44 and 25.48 UT. A CCD image taken by Tabur with a 0.28-m reflector yields the following precise position: α = 17^{h}47^{m}21^{s}.74, δ = −23^{o}28′23′′.1 (equinox 2000.0). W. Liller, Viña del Mar, Chile, provides the following position end figures from a CCD image taken on Oct. 27.1: 21^{s}.63 ± 0^{s}.05, 21′′.4 ± 0′′.4. Though the surrounding star field is crowded, Tabur adds that nothing appears at this location on Digitized Sky Survey images from 1991 (red) and 1987. Also, nothing appears at this location on previous images taken by Tabur on numerous dates back to Aug. 4, the most recent being Oct. 19. Magnitude estimates, visual unless otherwise noted: Oct. 21.1, [11.5 (Liller; red photograph); 26.354, 10.3 (A. Jones, Nelson, New Zealand); 26.42, 10.5 (Tabur); 26.42, 9.6 (Tabur; unfiltered CCD); 27.015, 9.98 (Liller; CCD, broadband V). F. Marang, South African Astronomical Observatory, reports the following infrared magnitudes: Oct. 25.75, H = 7.86, J = 8.36, K = 7.58; 26.74, 7.31, 7.88, 7.07.

SUPERNOVA 2002gw IN NGC 922 M. Hamuy, Carnegie Observatories, communicates: “S. Shectman and

I. Thompson, Carnegie Observatories, report that a spectrum (range 370– 540 nm) of SN 2002gw (cf. IAUC 7995), obtained on Oct. 24.25 UT with the Las Campanas Clay 6.5-m telescope (+ Boller & Chivens spectrograph), shows it to be a type-II supernova with prominent Balmer lines exhibiting P-Cyg profiles. I obtained a spectrum with the Baade 6.5-m telescope (+ LDSS2) on Oct. 25.20 that confirms the previous classification. The absorption component of the Hα line yields an expansion velocity of 9300 km/s, based on a redshift of 3100 km/s derived from the narrow Balmer emission lines of the host galaxy.”

POSSIBLE NOVA IN NGC 205 H. Yamaoka, Kyushu University; and T. Kato, Kyoto University, re-

port additional CCD magnitudes for this object (cf. IAUC 7984): Oct. 4.630 UT, ∼ 17.0 (Y. Mashitani, Tenkawa, Nara, Japan; prediscovery; via N. Fukuhara, Hyogo); 7.778, R = 18.5 (Z. Kereszty, Gyorujbarat, Hungary); 8.552, 18.1 (Y. Sano, Hokkaido, Japan). Yamaoka and Kato suggest it might be in a pre-maximum halt phase, exceptional for a very fast nova.

2002 October 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green